Tommy is standing 250 feet from the base of a building. The angle of elevation from his location to the top of the building is 50 degree. Based on the trigonometric function below, what is the height of the building, to the nearest foot?
Answer by richwmiller(17219)

opp/adj=x/250= tan50=1.1917535925942099587053080718604
x/250=1.1917535925942099587053080718604
x = 297.9383981485524896763270179651
x=298